Operations Report – April 22nd

rew 211 Operations Report 22-APR-2019

SOL: 1

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 11:30

On (last night): 7:30 pm

Off (this morning): 7 am

Solar— SOC 88 % at 6:45 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~420 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 115 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 6 units

Toilet tank emptied: yes

Deimos rover used: no (rover is out of order according to Director)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 88.8h

Beginning charge: 100

Ending charge: 50

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: yes

Hours: 89.8h

Beginning charge: 100

Ending charge: 83

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: We cleaned all helmets and boiled earphones. A new plastic box was cleaned too and now contains all suit-related supplies.

Summary of Hab operations: Floor of the toilets and the whole lower deck were cleaned, and the mess that was lying on the desk was sorted.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Two operations were done: building stands with wood and the handsaw for the UWB anchors by Simon, and Carl assembled its probes with clamps.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Someone slightly twisted a bit his ankle while climbing down a ladder on the upper deck. Not really a problem.

Questions, concerns and requests to Mission Support:

* The hot water heater does not seem to work. Water does flow into the hot water pipes and taps, but keeps being cold.

Perhaps a problem with the propane supply ?

* The pump that flows from the loft tank to the taps seems to be leaking. The cardboard on the ceiling of the rightmost room is getting wet.

It looks that at least one previous crew was aware as we found a few sponges placed under the pump.

We fear that some long-term damage could happen to the cardboard and the wood that is immediately underneath it.

Additionally there is electric wiring up there, and water could enter in contact with it.

* I doubt that we will have enough water for the whole mission. We are trying to limit our consumption, but as a crew of eight even drinking water represents quite a cost.